# Decentralized Aerial Manipulation of a Cable-Suspended Load using Multi-Agent Reinforcement Learning

### v1

@@ 3 Methods
Where $N$ is the total number of MAVs. The local policies have an observation space that only includes the load pose, relative goal terms, their own respective MAV state, and a one-hot vector $\bm{\mathit{e}}_{i}$ indicating their identity to enable role differentiation among homogeneous agents, as the policy network parameters are shared across all MAVs. The observation space for the $i$-th MAV is described as:
@@
$r_{t}=r_{t}^{\mathrm{pos}}+r_{t}^{\mathrm{ori}}+r_{t}^{\mathrm{down}}+r_{t}^{\mathrm{act}}+r_{t}^{\mathrm{br}}+r_{t}^{\mathrm{thrust}},$ (3)
@@
Action space and low-level controller To balance reliable sim-to-real transfer with sufficient control authority, the choice of action space is critical. Prior work in single MAV control demonstrates that high-level outputs (e.g., position or velocity) enhance robustness to disturbances and sim-to-real gaps but limit performance, whereas low-level outputs (e.g. snap) improve tracking precision at the cost of larger transfer discrepancies [41, 42]. To address this trade-off, we propose a mid-level action space in desired accelerations and body rates (ACCBR). This approach preserves adequate control capability while also being robust against uncertain disturbances and model mismatches from the cable-suspended load.
@@
$\bm{\mathit{z}}_{i,\mathrm{des}}=\frac{\bm{a}_{i,\mathrm{ref}}-\bm{g}-\bm{f}_{i,\mathrm{ext}}/m_{i}}{\|\bm{a}_{i,\mathrm{ref}}-\bm{g}-\bm{f}_{i,\mathrm{ext}}/m_{i}\|},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \bm{f}_{i,\mathrm{ext}}=\mathit{m}_{i}\bm{a}_{i,\mathrm{filtered}}-\bm{f}_{i,\mathrm{filtered}}$ (4)
@@
Training setup We train our method completely in simulation and achieve zero-shot transfer to real-world experiments. The simulation environment is built using NVIDIA’s Isaac Lab [43], and the MARL algorithms are modified from [44]. Training was conducted on a consumer-grade RTX 3090 GPU and completed in 17 hours. The network architecture is a 4-layer MLP of size [1024,512,256,128][1024,512,256,128][ 1024 , 512 , 256 , 128 ] for both the shared policies and the centralized critic. The inputs to the network are normalized stacked observation histories with history size $H=3$. For a complete overview of training details, network and agent parameters, we refer the readers to Appendix A.8.
The MAVs with the cable-suspended load spawn uniformly in a random location between 1-1- 1 and 111 in the xy-plane, between 0.50.50.5 and 1.51.51.5 along the z-axis, with a random heading orientation. The goal position is uniformly sampled from the same set, but also allows for pitch and roll angles of ±45$\pm$ 45± 45 degrees. Despite sampling of the goal is limited to the predefined sets, the policy is still able to generalize and reach goal poses outside of it during execution.
@@ 4.1 Real-world experiments
Setpoint tracking Our real-world experiments demonstrate agile pose control of three MAVs with a cable-suspended load, tracking a 2 m displacement with (30°, -20°, -90°) attitude commands. We compare our decentralized method with the state-of-the-art centralized NMPC approach [6] in Figure 3. Despite being fully decentralized, our method achieves comparable tracking performance with positional and attitude RMSEs of 0.52 m (vs 0.45 m) and 22.93 degrees (vs 16.24 degrees), respectively. Note that RMSE comparisons favor NMPC as it tracks a reference trajectory while we only track target poses, resulting in a larger RMSE in the transient area of the step command. We also show successful pose control with 4 MAVs (without cable slack), resulting in tracking RMSEs of 0.92 m and 42.67 degrees. The increased error, compared to the 3 MAV case, may be due to the system becoming overconstrained, which introduces more complex coordination and (cable) dynamics [18]. In terms of computational efficiency, we run the NMPC and our method onboard a Raspberry Pi 5 (2.4 GHz quad-core ARM Cortex-A76). Our method inferences in 6 ms at 100 Hz, versus NMPC’s 78 ms at 10 Hz. Crucially, while NMPC’s computation time grows exponentially with agent count, e.g. 174 ms and 267 ms for 5 and 6 agents respectively, our agent-independent approach maintains constant computation time regardless of team size.
Robustness against load model mismatch To evaluate robustness, we introduce five additional objects (0.216 kg, 15.4% of load mass), including four freely movable items that dynamically perturb both mass distribution and center of mass. Despite no inertia randomization during training, the system maintains strong tracking performance (0.63 m vs 0.60 m position RMSE; 26.93 degrees vs 26.49 degrees attitude error). The low-level feedback controller automatically compensates for these disturbances, demonstrating inherent robustness to model uncertainties. Experimental results are shown in Figure 4B.
@@
Heterogeneous agents Although our policy is trained under the assumption of homogeneous agents, it remains effective when deployed with heterogeneous agents. In this experiment, we let the load hover at a fixed point. Then we hacked one of the MAVs by replacing its RL policy with a model-based controller [11], and provided it with different setpoints to observe the behavior of the other two MAVs controlled by the RL policy. Specifically, we commanded the hacked MAV to move outwards on the y-axis by 0.7 m to pull the load away from the reference; we then commanded the hacked MAV to move inwards by 0.3 m to push it closer to the other two MAVs. Figure 4A provides a snapshot of the experiments. Since the policy is conditioned solely on the load pose and not on the states of the other agents, the two remaining MAVs utilizing the policy are able to compensate for load pose deviations from the reference. In contrast, the fully observable policy fails in these conditions because it depends on access to the states of all agents. Time series can be found in Appendix A.2.
In-flight failure of one MAV The effectiveness of our method with a heterogeneous agent setup and robustness against load model uncertainties also offers strong fault tolerance in the case of agent failure. In this experiment, we deliberately turned off the hacked MAV (one of the two on the same side). As a result, the load was controlled by the remaining two MAVs. Note that with only two MAVs, the load orientation around the line joining the remaining two attachment points becomes unactuated. Even worse, the failed MAV hangs underneath the load, leading to additional disturbances to the post-failure system. Despite that, our method allows the other two MAVs to effectively control the remaining 5 DoFs of the load. We show that the system is still able to yaw by -180 degrees and is also able to maintain position control by flying 0.5 meters down along the z-axis and maneuvering along the y-axis by 1 meter. The tracking results and snapshots of the setup after the failure are seen in Figure 4C. As in the heterogeneous agent case, the remaining agents can compensate for the missing MAV since the policy operates independently of other agents’ states, thereby avoiding unstable behavior in out-of-distribution scenarios. In contrast, the fully observable policy fails under these conditions due to its reliance on the states of all agents. Time series illustrating both scenarios can be found in Appendix A.3.
@@ 4.2 Comparison among different action and observation spaces
We compare our selected observation and action space with other alternatives. The comparisons are performed in a simulation environment for safety reasons. We ran the Agilicious flight stack together with the Gazebo simulator [45] with quadrotor and sensor plugins provided by the RotorS [46] library, which introduces sensor noise, aerodynamic disturbances, and potential system latencies in a ROS environment. All policies are trained on a limited budget of 1 billion environment steps (10 hours of training), and are evaluated 10 times in the Gazebo environment.
@@
Figure 5: Positional and attitude errors comparing different action spaces at test time in the Gazebo environment.
@@
Action space Pos RMSE Att RMSE ACCBR 0.64 $\pm$ 0.00 33.87 $\pm$ 0.91 CTBR NaN NaN ACC 0.54 $\pm$ 0.00 87.89 $\pm$ 1.85 VEL 0.56 $\pm$ 0.06 25.74 $\pm$ 1.49 Not able to take off
Figure 6: Pose tracking RMSEs of different action spaces at test time in the Gazebo environment.
@@
However, while VEL yields superior RMSE, Figure 6 reveals that it induces hazardous oscillations during execution. In contrast, ACCBR exhibits more stable hovering despite higher initial errors. For real-world tasks like inspection or delivery—where stability is critical—we argue that ACCBR is the safer and preferred choice.
@@ 6 Limitations
Our method requires pose measurement of the load, which is not often practical beyond lab environments. In our experiment, we require an external motion capture system to provide high-frequency load pose measurement. For future real-world outdoor deployment, onboard sensing (e.g., a downward-facing camera for load pose estimation and SLAM for MAV localization) would be necessary. This would introduce new challenges, such as different reference frames for the load and MAVs—requiring additional transformation and synchronization—, observation delays, imperfect state estimates, and sensor noise. Additionally, our current framework does not address obstacle avoidance, as we assume collision-free paths to the goal—an unrealistic assumption in unstructured environments. Future work will focus on integrating a robust perception stack and obstacle avoidance capabilities.
@@ Acknowledgments
The authors would like to thank Dr. Yunlong Song, Dennis Benders and Shlok Deshmukh for the insightful discussions, and Maurits Pfaff and Kseniia Khomenko for their help with the experiments.
Appendix A Supplementary Materials
@@ A.6 Performance with different history lengths
Figure 13: Training curves comparing different history lengths for the partially observable policy.
Figure 14: Mean reward of policies with different history lengths over 10 runs at test time in the Gazebo environment.
Figures 14 and 14 show that including historical observations has little impact on performance. We hypothesize that the load’s pose—even without historical data—contains enough information to estimate the other agents’ states, enabling implicit communication among the MAVs. Further investigation into the role of history in more complex scenarios, such as those with higher noise or additional MAVs, is left for future work.
@@ A.8 Training configuration
[…]
@@
  • any of the cable tensions are below a specified threshold. (>3>3> 3 MAVs)

### v2

@@ 3 Methods
Where $N$ is the total number of MAVs. The local policies’ observation space only includes the load pose, relative goal terms, their own respective MAV state, and a one-hot vector $\bm{\mathit{e}}_{i}$ indicating their identity to enable role differentiation among homogeneous agents, as the policy network parameters are shared across all MAVs. The observation space for the $i$-th MAV is described as:
@@
$r_{t}=r_{t}^{\mathrm{pos}}+r_{t}^{\mathrm{ori}}+r_{t}^{\mathrm{down}}+r_{t}^{\mathrm{act}}+r_{t}^{\mathrm{br}}+r_{t}^{\mathrm{thrust}}$ (3)
@@
Action space and low-level controller To balance reliable sim-to-real transfer with sufficient control authority, the choice of action space is critical. Prior work in single MAV control demonstrates that high-level outputs (e.g., position or velocity) enhance robustness to disturbances and sim-to-real gaps but limit performance, whereas low-level outputs (e.g., snap) improve tracking precision at the cost of larger transfer discrepancies [41, 42]. To address this trade-off, we propose a mid-level action space in desired accelerations and body rates (ACCBR). This approach preserves adequate control capability while also being robust against uncertain disturbances and model mismatches from the cable-suspended load.
@@
$\bm{\mathit{z}}_{i,\mathrm{des}}=\frac{\bm{a}_{i,\mathrm{ref}}-\bm{g}-\bm{f}_{i,\mathrm{ext}}/m_{i}}{\|\bm{a}_{i,\mathrm{ref}}-\bm{g}-\bm{f}_{i,\mathrm{ext}}/m_{i}\|},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \bm{f}_{i,\mathrm{ext}}=\mathit{m}_{i}\bm{a}_{i,\mathrm{filtered}}-\bm{f}_{i,\mathrm{filtered}}$ (4)
@@
Training setup We train our method completely in simulation and achieve zero-shot transfer to real-world experiments. The simulation environment is built using NVIDIA’s Isaac Lab [43], and the MARL algorithms are modified from [44]. Training was conducted on a consumer-grade RTX 3090 GPU and completed in 17 hours. The network architecture is a 4-layer MLP of size $[1024,512,256,128]$ for both the shared policies and the centralized critic. The inputs to the network are normalized stacked observation histories with history size $H=3$. For a complete overview of training details, network and agent parameters, we refer the readers to Appendix A.8.
The MAVs with a cable-suspended load spawn uniformly between 1-1 and $1$ in $xy$, $0.5$ and $1.5$ in $z$, with a random heading. The goal position is sampled from the same range, but also allows pitch and roll of $\pm 45^{\circ}$. Despite sampling of the goal is limited to the predefined sets, the policy is still able to generalize and reach goal poses outside of it during execution.
@@ 4.1 Real-world experiments
Setpoint tracking Our real-world experiments demonstrate agile pose control of three MAVs with a cable-suspended load, tracking a 2 m displacement with (30°, -20°, -90°) attitude commands. We compare our decentralized method with the state-of-the-art centralized NMPC approach [6] in Figure 3. Despite being fully decentralized, our method achieves comparable tracking performance with positional and attitude RMSEs of 0.52 m (vs 0.45 m) and 22.93° (vs 16.24°), respectively. Note that RMSE comparisons favor NMPC as it tracks a reference trajectory while we only track target poses, resulting in a larger RMSE in the transient area of the step command. The time-to-target (error $\text{0.10}\,\text{m}/\text{10}^{\circ}$) is 6.84 s for NMPC vs. 8.36 s for ours, and the final displacement (RMSE) is 0.05 m /4.02 for NMPC vs. 0.04 m /5.78 for ours. We also show successful pose control with 4 MAVs (without cable slack), resulting in tracking RMSEs of 0.92 m and 42.67°. The increased error, compared to the 3 MAV case, may be due to the system becoming overconstrained, which introduces more complex coordination and (cable) dynamics [18]. In terms of computational efficiency, we run the NMPC and our method onboard a Raspberry Pi 5 (2.4 GHz quad-core ARM Cortex-A76). Our method inferences in 6 ms at 100 Hz, versus NMPC’s 78 ms at 10 Hz. Crucially, while NMPC’s computation time grows exponentially with agent count, e.g., 174 ms and 267 ms for 5 and 6 agents respectively, our agent-independent approach maintains constant computation time regardless of team size.
Robustness against load model mismatch To evaluate robustness, we add objects (0.216 kg, 15.4% of load mass) to the load, including four freely movable items that dynamically perturb both mass distribution and center of mass. Despite no inertia randomization during training, the system maintains strong tracking performance (0.63 m vs 0.60 m position RMSE; 26.93° vs 26.49° attitude RMSE. The low-level feedback controller automatically compensates for these disturbances, demonstrating inherent robustness to model uncertainties. Experimental results are shown in Figure 4B.
@@
Heterogeneous agents Although our policy is trained under the assumption of homogeneous agents, it remains effective when deployed with heterogeneous agents. In this experiment, we let the load hover at a fixed point. Then we hacked one of the MAVs by replacing its RL policy with a model-based controller [11], and provided it with different setpoints to observe the behavior of the other two MAVs controlled by the RL policy. Specifically, we commanded the hacked MAV to move outwards on the y-axis by 0.7 m to pull the load away from the reference; we then commanded the hacked MAV to move inwards by 0.3 m to push it closer to the other two MAVs. Figure 4A provides a snapshot of the experiments. Since the policy is conditioned solely on the load pose and not on the states of the other agents, the two remaining MAVs utilizing the policy can compensate for load pose deviations from the reference. In contrast, the fully observable policy fails in these conditions due to the dependence on the states of all agents. Time series can be found in Appendix A.2.
In-flight failure of one MAV The effectiveness of our method with a heterogeneous agent setup and robustness against load model uncertainties also offers strong fault tolerance in the case of agent failure. In this experiment, we deliberately turned off the hacked MAV (one of the two on the same side). As a result, the load was controlled by the remaining two MAVs. Note that with only two MAVs, the load orientation around the line joining the remaining two attachment points becomes unactuated. Even worse, the failed MAV hangs underneath the load, leading to additional disturbances to the post-failure system. Despite that, our method allows the other two MAVs to effectively control the remaining 5 DoFs of the load. We show that the system is still able to yaw by -180° and is also able to maintain position control by flying 0.5 meters down along the z-axis and maneuvering along the y-axis by 1 meter. The tracking results and snapshots of the setup after the failure are seen in Figure 4C. As in the heterogeneous agent case, the remaining agents can compensate for the missing MAV since the policy operates independently of other agents’ states, thereby avoiding unstable behavior in out-of-distribution scenarios. In contrast, the fully observable policy fails under these conditions due to its reliance on the states of all agents. Time series illustrating both scenarios can be found in Appendix A.3.
@@ 4.2 Comparison among different action and observation spaces
We compare our selected observation and action spaces with alternatives in simulation for safety. The Agilicious flight stack is used with the Gazebo simulator [45] and RotorS [46] plugins, which add sensor noise, aerodynamic disturbances, and system latencies in a ROS environment. All policies are trained for 1 billion environment steps (10 h) and evaluated 10 times in Gazebo.
@@
[…]
@@
However, while VEL achieves lower RMSE, Figure 6 shows it causes hazardous oscillations. ACCBR offers more stable hovering despite higher initial errors, making it safer and preferable for stability-critical tasks like inspection or delivery.
@@ 6 Limitations
Our method requires pose measurement of the load, which is not often practical beyond lab environments. In our experiment, we require an external motion capture system to provide high-frequency load pose measurement. For future real-world outdoor deployment, onboard sensing (e.g., a downward-facing camera for load pose estimation and SLAM for MAV localization) would be necessary. This would introduce new challenges, such as observation delays, imperfect state estimates, sensor noise, and different reference frames for the load and MAVs that require alignment and synchronization. Additionally, our current framework does not address obstacle avoidance, as we assume collision-free paths to the goal—an unrealistic assumption in unstructured environments. Future work will focus on integrating a robust perception stack and obstacle avoidance capabilities.
@@ Acknowledgments
The authors would like to thank Dr. Yunlong Song, Dr. Dennis Benders, and Shlok Deshmukh for the insightful discussions, and Maurits Pfaff and Kseniia Khomenko for their help with the experiments. This work is funded by the Dutch Research Council (NWO) under Grant 20256 for the project "Accurate Aerial Manipulation under Uncertainties" and by the European Union under the ERC grant INTERACT, 101041863. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Appendix
@@ A.6 Performance with different history lengths
[…]
Figures 16 and 16 show that including historical observations has little impact on performance. We hypothesize that the load’s pose—even without historical data—contains enough information to estimate the other agents’ states, enabling implicit communication among the MAVs. Further investigation into the role of history in more complex scenarios, such as those with higher noise or additional MAVs, is left for future work.
@@ A.8 Training configuration
The inputs to the network are normalized stacked observation histories with history size $H=3$. We also implement a form of advantage filtering [51] where $50\%$ of the samples with the lowest advantage magnitude are dropped. This approach prioritizes learning from the most informative state transitions—specifically the underexplored extremes of the data distribution where actions have a clearly better or worse outcome—thereby improving data efficiency during training. For a complete overview of the network and agent parameters, we refer the readers to Table 5.
@@
  • any of the cable tensions are below a specified threshold. (>3>3 MAVs)